OpenStudy (anonymous):

7a^-20=0 quadratic formula how do you solve

OpenStudy (anonymous):

not clear what you wrote is it \[7a^2-20=0\]?

OpenStudy (anonymous):

\[7a^2=20 ......> a^2=20/7.......> a= \pm \sqrt{20/7}\]